OpenStudy (erod0):

Rewrite the logarithmic expression as a single logarithm and simplify the result.

OpenStudy (erod0):

\[\ln(\sin \theta^2)-\ln(\tan \theta^2)\]

OpenStudy (solomonzelman):

\(\color{black}{\displaystyle \ln\left[(\sin \theta)^2\right]-\ln\left[(\tan \theta)^2\right]}\) ?

OpenStudy (solomonzelman):

or, is it for some angle \(\theta^2\) ?

OpenStudy (erod0):

the first one

OpenStudy (solomonzelman):

Another correct way to write the first one is: \(\color{black}{\displaystyle \ln\left[\sin^2\theta\right]-\ln\left[\tan^2\theta\right]}\)

OpenStudy (solomonzelman):

anyway ...

OpenStudy (erod0):

ok

OpenStudy (solomonzelman):

The first rule you need is: 1. \(\color{black}{\displaystyle \ln\left[a^k\right]=k\times \ln[a]}\)

OpenStudy (solomonzelman):

Apply this to \(\color{black}{\displaystyle \ln\left[(\sin \theta)^2\right]}\) and to \(\color{black}{\displaystyle \ln\left[(\tan \theta)^2\right]}\). Then write what you get.

OpenStudy (solomonzelman):

(Don't be afraid to get anything wrong if you want to attempt.)

OpenStudy (erod0):

ok gimme a minute

OpenStudy (erod0):

2 ln |cos\[\theta\]

OpenStudy (erod0):